### v1

# arXiv:2603.00918v1[cs.CV] 01 Mar 2026

Abstract
Text-to-image generation powers content creation across design, media, and data augmentation. Post-training of text-to-image generative models is a promising path to better match human preferences, factuality, and improved aesthetics. We introduce ARC (Adaptive Rewarding by Self-Confidence), a post-training framework that replaces external reward supervision with an internal self-confidence signal, obtained by evaluating how accurately the model recovers injected noise under self-denoising probes. ARC converts this intrinsic signal into scalar rewards, enabling fully unsupervised optimization without additional datasets, annotators, or reward models. Empirically, by reinforcing high-confidence generations, ARC delivers consistent gains in compositional generation, text rendering and text-image alignment over the baseline. We also find that integrating ARC with external rewards results in a complementary improvement, with alleviated reward hacking.
Figure 1: Qualitative examples of ARC on Pick-a-Pic dataset [30]. Best viewed on electronics.
1 Introduction
Text-to-image (T2I) generation has advanced rapidly with the rise of diffusion and flow-based models, delivering high-fidelity, diverse images from natural language prompts [55, 56, 48, 46, 10, 9, 16, 22]. These models now support a broad range of applications: controllable image editing and inpainting [7, 6, 61, 3, 60, 76, 87]; serving as powerful priors or pre-trained components for text-to-video diffusion models [92, 71, 24, 26, 31, 70]; data creation and augmentation pipelines for downstream perception tasks [64, 80, 73]; and text-to-3D (and 4D) reconstruction via score distillation sampling [50, 62, 74, 65, 1]. Recent studies show that post-training text-to-image generative models via reinforcement learning can tailor these models to show dramatic improvements in visual appeal and aesthetic quality [69, 5, 36], typically by optimizing external rewards derived from human preference models [30, 75, 78] or task-specific validators [21, 14].
However, defining a scalable and reliable reward for “good” images remains challenging [30, 75, 78, 33, 67]. There are numerous, weakly-aligned criteria a good image has to satisfy, e.g., compositionality, text rendering, aesthetics, and text–image alignment, whose relative importance shifts across domains and prompts [33]. In practice, external-reward post-training is also vulnerable to over-optimization: optimizing a narrow critic can induce reward hacking and regressions on non-target capabilities, degrading coverage or faithfulness even as the targeted score rises [69, 5, 36]. Human-preference based reward models [30, 78, 72] are popular for their efficacy, but require large-scale annotation for training. Operationally, external rewards require running additional evaluators (preference/OCR/safety models) alongside the generator during training, increasing pipeline complexity.
Despite extensive progress in extrinsically supervised post-training, relying on intrinsic signals remains under-explored for text-to-image generation. In this work, we aim to answer a fundamental question: can internal feedback from the text-to-image generator itself provide meaningful signals for post-training? To this end, we introduce Adaptive Rewarding by Self-Confidence (ARC), a post-training framework that relies on intrinsic self-certainty as the model-native reward. Inspired by Score Distillation Sampling [50, 62, 74], which leverages a pretrained text-to-image generator as a critic for text-to-3D or -4D generation, we propose to leverage a text-to-image generator to critique its own generation. Concretely, given a sampled latent $z_{0}$, we re-noise $z_{0}$ to selected timesteps $t\in\mathcal{T}$ using the model’s forward noising schedule, and measure how well the model reconstructs the injected noise, rewarding small residuals as ’high confidence’. Our hypothesis is that large-scale pretraining in diffusion generative models endows them with priors over real images and text–image alignment, so the model’s self-confidence would align strongly with text-alignment and realism.
Empircally, ARC yields consistent gains in compositional generation [21], text rendering [14], and text–image alignment [53], while modestly improving human-preference scores [30, 75, 78, 72] without external rewards. Qualitative side-by-side comparisons and a comprehensive user study corroborate these trends. Together, the results indicate that intrinsic self-confidence aligns with key facets of image generation. Moreover, applying ARC on top of an extrinsically post-trained model (i.e., using external rewards) recovers additional improvements in compositionality, text rendering, and alignment, with only slight drops on the targeted external metric, evidencing that intrinsic and extrinsic rewards are complementarily beneficial for post-training stronger T2I generators.
The key contributions of our work are as follows:
  • We present ARC (Adaptive Rewarding by self-Confidence), a post-training framework to use self-confidence as a reward for text-to-image generators.
  • We define ARC’s self-confidence reward by re-noising the model’s own outputs and rewarding accurate recovery of the injected noise - a training-aligned signal that serves as a principled self-confidence score.
  • Across standard benchmarks and a comprehensive user study, ARC yields consistent gains in compositionality, text rendering, and text–image alignment, while modestly improving human-preference metrics.
  • ARC complements external-reward pipelines: applying ARC on top of externally post-trained models improves non-target capabilities (compositionality, text rendering, alignment) with only mild trade-offs on the targeted external metric, mitigating reward hacking.
2 Related Work
Text-to-image generative models. Text-to-image generation is a rapidly advancing field, which was initially dominated by diffusion models [2, 10, 9, 46, 48, 55, 56]. Recent studies tend to focus more on flow matching [16, 3, 63] and sequence models [81, 8, 42, 68] for improved efficiency in train/test time, and enhanced image generation performance. Various innovative techniques are being proposed across architectures [46, 16, 3], image recaptioning [4, 10, 9] and tokenization [68, 82, 29] to further improve the potential of text-to-image generation. In this work, we focus on reinforcement-learning based post-training of text-to-image models to improve a pretrained text-to-image generation model, by using the self-confidence of the generative model as the intrinsic reward.
Text-to-image model alignment via post-training. Post-training is emerging as an effective paradigm to align existing text-to-image models to output image to desired directions e.g., human preference. This post-training can take the form of direct fine-tuning given differentiable rewards [51, 13, 78, 52] or Reward Weighted Regression (RWR) [47, 17, 32, 15]. Some post-training schemes build on reinforcement learning to leverage PPO [58]-style policy gradients [5, 18, 44, 25, 89], or perform Direct Preference Optimization (DPO) or its variants [54, 69, 83, 38, 79, 86, 19, 34, 39, 34]. More recently, Flow-GRPO [37] introduces GRPO [59] for flow matching models, by converting the ODE of flow matching sampling to SDEs to inject stochasticity. However, leveraging external rewards results in enhanced costs from having another model during training and also increased risks of reward hacking [54, 69, 39, 38]. In this work, we formulate the self-certainty of text-to-image generated outputs as the ability to recover the noise injected to its own generated outputs, which we leverage as the intrinsic signal to perform post-training, resulting in improved text-to-image generation without reward hacking.
Intrinsic signals for post-training. Intrinsic signals for post-training have recently gained traction in language modeling as scalable alternatives to human-labeled preference data, leveraging self-derived feedback such as confidence/uncertainty estimates, self-evaluation, and self-consistency to guide reinforcement learning or preference optimization without annotators [85, 11, 84, 49, 12, 90, 77, 93, 88]. Recently, Inituitor [90] showed that using self-certainty as a confidence-based intrinsic reward enables single-agent reinforcement learning across diverse tasks without relying on explicit feedback, gold labels, or environment-based validation. Bringing the same principle to text-to-image generation is non-trivial: generation proceeds along continuous denoising trajectories and likelihoods are implicit, unlike token-level discrete objectives in LLMs. In this work, we instantiate self-certainty of flow-matching models as their ability to recover the noise injected to its generated outputs, inspired by score-distillation sampling semantics [50, 62]. This enables us to obtain dense, on-policy feedback without additional datasets or reward models. Empirically, we show that this confidence signal aligns with key facets of image generation, i.e., compositionality, text rendering, and text–image alignment.
3 Preliminary: GRPO for Flow Matching
3.1 Flow Matching and Rectified Flow
Flow matching bypasses score learning in conventional diffusion models [27, 66] by directly regressing the target velocity of a transport ODE along a user-chosen path between data and a reference distribution [35, 40]. Recent state-of-the-art generative models [16, 3, 70, 31] adopt the Rectified Flow (RF) framework. Specifially, let $x_{0}\!\sim p_{\text{data}}$ and $x_{1}\!\sim p_{1}$ (e.g., $\mathcal{N}(0,I)$); RF chooses the straight-line path
$x_{t}=(1-t)\,x_{0}+t\,x_{1},$ (1)
for which the target velocity is constant in $t$:
$v^{\star}=\partial_{t}x_{t}=x_{1}-x_{0}.$ (2)
Training reduces to direct regression of this constant velocity at random $(x_{t},t)$ pairs:
$\mathcal{L}(\theta)=\mathbb{E}_{x_{0}\sim p_{\text{data}},\,x_{1}\sim p_{1},\,t\sim\mathcal{U}[0,1]}\Big\|v-v_{\theta}(x_{t},t)\Big\|_{2}^{2}.$ (3)
After training, sampling solves the deterministic ODE
$\frac{\mathrm{d}x_{t}}{\mathrm{d}t}=v_{\theta}(x_{t},t),\qquad t:1\!\to\!0,$ (4)
starting from $x_{1}\!\sim p_{1}$ and transporting to $x_{0}$.
Figure 2: Overview of ARC. Given a text prompt $c$, we generate $G$ different latents. Without decoding, we re-noise the latents using $K$ noise probes across $t\in\mathcal{T}\subset[0,1]$. For each generated latent $z_{0}^{(i)}$, we formulate the text-to-image generative model’s self-confidence of the generated latent as the ability to denoise the re-noised latent. We leverage this self-confidence as an internal reward scalar value, which we use to post-train the text-to-image generative model using GRPO [59, 37]. We omit the KL term in this figure for better readability.
3.2 GRPO for Flow Matching
For a policy $\pi_{\theta}$, we consider a policy-gradient objective that maximizes expected cumulative reward while regularizing updates toward a reference policy $\pi_{\mathrm{ref}}$ via a KL penalty:
$\max_{\theta}\mathbb{E}_{(s_{0},a_{0},\ldots,s_{T},a_{T})\sim\pi_{\theta}}$ (5)
$\Big[\sum_{t=0}^{T}R(s_{t},a_{t})-\beta\sum_{t=0}^{T}D_{\mathrm{KL}}\!\left(\pi_{\theta}(\cdot\mid s_{t})\,\|\,\pi_{\mathrm{ref}}(\cdot\mid s_{t})\right)\Big],$
where $R(s_{t},a_{t})$ is the per-step reward. Group Relative Policy Optimization (GRPO) [59] proposes to use a group relative formulation to estimate the advantage for each sample to optimize Eq. 5.
Flow-GRPO [59] proposes to integrate GRPO into flow matching models for online RL post-training. The iterative denoising process in flow matching can be formulated as a Markov Decision Process [5]: given a text prompt $c$, the flow model $p_{\theta}$ samples a group of $G$ images $\{x^{i}_{0}\}_{i=1}^{G}$ and the corresponding sampling trajectories $\{(x_{T}^{i},x_{T-1}^{i},\cdots,x_{0}^{i})\}_{i=1}^{G}$. The advantage of the $i$-th image is calculated by normalizing the group-level rewards:
$\hat{A}^{i}_{t}=\frac{R(x_{0}^{i},c)-\textrm{mean}(\{R(x_{0}^{i},c)\}_{i=1}^{G}\})}{\textrm{std}(\{R(x_{0}^{i},c)\}_{i=1}^{G})}$ (6)
Finally, GRPO optimizes the policy model by maximizing $\mathcal{J}_{\textrm{Flow-GRPO}}=\mathbb{E}_{c\sim C,\{x^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\textrm{old}}}(\cdot|c)}f(r,\hat{A},\theta,\epsilon,\beta)$, where
$f(r,\widehat{A},\theta,\epsilon,\beta)$ $=\underset{i,t}{\mathrm{mean}}\Big[\min\!\big(r_{t}^{\,i},\,\mathrm{clip}_{\epsilon}(r_{t}^{\,i})\big)\,\widehat{A}_{t}^{\,i}\Big]\;-\;\beta\,\overline{D}_{\mathrm{KL}},$ (7)
$\overline{D}_{\mathrm{KL}}$ $=\underset{t}{\mathrm{mean}}\;D_{\mathrm{KL}}\!\left(\pi_{\theta}(\cdot\mid s_{t})\,\|\,\pi_{\mathrm{ref}}(\cdot\mid s_{t})\right),$
$\mathrm{clip}_{\epsilon}(r)$ $\triangleq\;\mathrm{clip}(r,1-\epsilon,1+\epsilon).$
and $r^{i}_{t}(\theta)=\frac{p_{\theta}(x^{i}_{t-1}|x^{i}_{T},c)}{p_{\theta_{\textrm{old}}}(x^{i}_{t-1}|x^{i}_{T},c)}$. Flow-GRPO then converts the deterministic ODE of Eq. 4 into an equivalent SDE that matches the original model’s marginal probability function at all timesteps, in order to meet the GRPO policy update requirements, e.g., stochasticity is necessary for increased randomness, which is directly related to “exploration” in RL post-training. We adopt Flow-GRPO in our work to post-train flow-matching based text-to-image generative models.
4 Adaptive Rewarding by Self-Confidence
Overview. We present ARC (Adaptive Rewarding by Self-Confidence), a post-training method for text-to-image generators that eschews reliance on external reward models. ARC leverages the model’s own self-confidence as an intrinsic reward: after generating an output, we re-noise it at selected timesteps and score how accurately the model recovers the injected noise. Aggregating these per-timestep recovery errors yields a single on-policy scalar reward for reinforcement learning. In the following, we detail the computation of the self-confidence reward (Sec. 4.1), and the stabilization and efficiency techniques used for using ARC (Sec. 4.2). An overview of ARC is shown in Fig. 2.
4.1 Intrinsic Self-Confidence Reward
Sampling group of images for GRPO. Given a text prompt $c$, we sample $G$ independent reverse trajectories in the latent space $\mathcal{Z}$ under the flow policy $\pi_{\theta}$:
$z_{T}^{(i)}\sim\mathcal{N}(0,I),z_{t-1}^{(i)}\sim\pi_{\theta}\!\left(\cdot\mid z_{t}^{(i)},\,c\right),i=1,\ldots,G.$ (8)
This produces terminal latents $\{z_{0}^{(i)}\}_{i=1}^{G}$ and trajectories $\{(z_{T}^{(i)},z_{T-1}^{(i)},\ldots,z_{0}^{(i)})\}_{i=1}^{G}$. Using multiple independent draws yields the group required for group-relative advantage normalization in GRPO. While we can sample $G$ different images from the same initial noise $z_{T}$ due to the added stochasticity from [37], we sample different initial noise for improved exploration for efficient GRPO training.
Sampling noise probes for re-noising. We draw a shared set of $K$ noise probes in latent space:
$\epsilon^{(m)}\sim\mathcal{N}(0,I),\qquad m=1,\ldots,K,$ (9)
so that candidate $i$ and candidate $j$ are perturbed by the same probes $\{\epsilon^{(m)}\}_{m=1}^{K}$. For rectified flow, we re-noise a terminal latent $z_{0}^{(i)}$ via the linear forward kernel
$z_{t}^{(i,m)}\;=\;(1-t)\,z_{0}^{(i)}\;+\;t\,\epsilon^{(m)},\qquad t\in\mathcal{T}\subset[0,1],$ (10)
where $\mathcal{T}$ is the set of re-noising levels used for evaluation. We take $K$ even ($K\!\geq\!2$) and use antithetic pairing to enforce exact mean zero within the probe set, i.e., $\epsilon^{(m+K/2)}=-\,\epsilon^{(m)}$ for $m=1,\ldots,K/2$.
Calculating self-confidence. For each noised latent $z_{t}^{(i,m)}$ (Eq. (10)), we query the flow-matching model’s velocity field $v_{\theta}(z_{t}^{(i,m)},t,c)$. Under the rectified-flow parameterization, the velocity predicts a linear transform of the injected noise; specifically, we recover a noise estimate via
$\widehat{\epsilon}_{\theta}\!\left(z_{t}^{(i,m)},t,c\right)\;=\;v_{\theta}\!\left(z_{t}^{(i,m)},t,c\right)\;+\;z_{0}^{(i)}.$ (11)
We then measure the reconstruction error against $\epsilon^{(m)}$:
$\mathrm{MSE}_{i,t}\;=\;\frac{1}{K}\sum_{m=1}^{K}\left\|\widehat{\epsilon}_{\theta}\!\left(z_{t}^{(i,m)},t,c\right)-\epsilon^{(m)}\right\|_{2}^{2}.$ (12)
To turn small errors into large rewards while stabilizing dynamic range, we use the negative log transform,
$S_{i,t}\;=\;-\,\log\!\big(\mathrm{MSE}_{i,t}+\delta\big),$ (13)
where $\delta>0$ avoids $\log 0$. This choice (i) approximates a Gaussian log-likelihood score under an i.i.d. noise model, (ii) compresses outliers, and (iii) yields additive contributions across timesteps. Aggregating over a set of re-noising levels $\mathcal{T}\subset[0,1]$ gives the scalar intrinsic reward
$R_{\mathrm{ARC}}\big(z_{0}^{(i)},c\big)\;=\;\frac{1}{\sum_{t\in\mathcal{T}}w(t)}\sum_{t\in\mathcal{T}}w(t)\,S_{i,t}.$ (14)
We use $w(t)=1$ in practice for simplicity. Note that external rewards typically operate in pixel space, $R_{\mathrm{ext}}(x^{(i)},c)$, where $x^{(i)}=\mathrm{Dec}(z_{0}^{(i)})$ for a fixed decoder $\mathrm{Dec}\!:\mathcal{Z}\!\to\!\mathcal{X}$. In contrast, $R_{\mathrm{ARC}}$ is computed directly in latent space, avoiding decoding and keeping the signal model-native.
4.2 Stabilization and Efficiency Techniques
Denoising reduction for efficient training. Following Flow-GRPO [37], we shorten the reverse-time horizon during training by subsampling the denoising steps. This reduces compute without degrading downstream gains: e.g., while SD3.5 uses $40$ sampling steps at inference, we use 10 denoising steps during training . We empirically find that this does not sacrifice image quality at test time, but facilitates substantially faster training.
Timestep selection for self-confidence probing. We probe self-confidence at the exact scheduler timesteps used by the SD3.5 sampler (same discretization and indices), ensuring the suffix window align with the generation trajectory. This on-policy, solver-aware choice avoids distributional mismatch between sampling and self-confidence probing, yielding more reliable credit assignment across steps.
Training on selective timesteps. We observed that using all training denoising trajectories easily leads to collapse (e.g., blank / textureless images), a form of reward hacking in which the model steers latents toward regimes that make injected noise overly easy to predict. We mitigate this by optimizing the sampled trajectories from only a suffix of the training schedule, i.e., a fixed percentage of the later reverse steps, where the denoising task remains informative but less exploitable. Let $\mathcal{T}_{\mathrm{train}}\!\subset\!\mathcal{T}$ denote this suffix window ($\lvert\mathcal{T}_{\mathrm{train}}\rvert=\lceil\rho\,\lvert\mathcal{T}\rvert\rceil$); we apply GRPO losses only on $t\!\in\!\mathcal{T}_{\mathrm{train}}$, which stabilizes learning without collapse.
Self-Confidence calculation with CFG. Although $G$ images are sampled with CFG for GRPO training, our ARC self-confidence is computed without CFG. CFG forms a mixture field $v_{\mathrm{cfg}}=v_{\mathrm{uncond}}+s\,(v_{\mathrm{cond}}-v_{\mathrm{uncond}})$; this mixture would calculate a self-certainty based on the guided proxy rather than the base conditional model trained on data. In our experiments, omitting CFG during self-confidence computation yields stronger and more stable improvements.
Online calculation of self-confidence. When computing self-confidence for ARC, we can either compute it (1) online, using the model being trained ($\pi_{\theta}$), or (2) offline, using a fixed based model ($\pi_{\textrm{ref}}$) for the self-confidence calculation. While we do not observe severe over-optimization [20] when computing the self-confidence in an offline manner, we notice that the performance improves when we compute self-confidence rewards in an online manner. We conjecture that as the model improves through ARC post-training, the stability and reliability of self-confidence improves as well, leading to a stronger improvement in performance.
Task-specific Image Quality Human Preference
Model GenEval OCR ClipScore Aesthetic PickScore HPSv2.1 ImageReward UnifiedReward
SD-XL 0.55 0.14 0.287 5.60 22.42 0.280 0.76 2.93
SD3.5-L 0.71 0.68 0.289 5.50 22.91 0.288 0.96 3.25
SD3.5-M 0.65 0.61 0.282 5.36 22.34 0.279 0.84 3.08
+ ARC (Ours) 0.71 0.67 0.288 5.39 22.41 0.278 0.87 3.11
[…]
Table 1: Quantitative results of ARC. We evaluate ARC post-training on base models of SD3.5 [16], on GenEval [21], Text Rendering, human preference reward models [30, 75, 78, 72], and image quality metrics. We show that ARC post-training yields consistent gains across text-to-image generative models on different quantitative metrics.
Figure 3: User study against baseline SD3.5-M [16] on PartiPrompts [56] and HPSv2 [75]. The user study shows that ARC post-training yields favorable visual realism/appeal, and text-image alignment.
5 Experiments
5.1 Implementation details
We use a group size $G=16$ and number of noise probes $K=8$ with antithetic pairing in our experiments. While we do not need external reward models for training, we still need text prompts to generate the terminal latents for training; we use the train set of the visual text rendering task [14] from Flow-GRPO [37], which holds longer and more informative prompts compared to Pick-a-Pic [30] or GenEval [21]. We use parameter-efficient post-training via leveraging LoRA [28, 43] with rank $r=32$ and scaling factor $\alpha=64$. We use the AdamW [41] optimizer with constant learning rate of 3e-4, and the KL regularizer weight $\beta=0.04$. In $\lvert\mathcal{T}_{\mathrm{train}}\rvert=\lceil\rho\,\lvert\mathcal{T}\rvert\rceil$, we set $\rho=0.6$, which we find to show improvements without reward hacking or training collapse. An image resolution of $\times512$ is used for both training and testing. We use CFG guidance of of 7.0 at inference. All experiments are carried out on $8\timesNVIDIA$ RTX PRO 6000 Blackwell GPUs. We include more training details in the supplementary materials.
5.2 Evaluation setting
(1) Compositional image generation. We evaluate on GenEval [21] that consists of complex compositional prompts including object counting, attribute bounding or spatial relations. Evaluation is performed across six tasks of position, counting, attribute binding, colors, two objects, and single object. We adhere to the official evaluation pipeline to detect object bounding boxes and colors, then infers their spatial relations, when given a generated image. The rewards are then calculated in a rule-based manner e.g. for object counting, $r=1-\frac{|N_{\textrm{gen}}-N_{\textrm{ref}}|}{N_{\textrm{ref}}}$, where $N_{\textrm{gen}}$ is the number of generated objects, while $N_{\textrm{ref}}$ is the specified number of objects in the prompt.
(2) Visual text rendering. We use the 1,000 GPT4o [45]-generated test prompts from [37]. In each prompt, the exact string that should appear in the image (i.e. target text) is specified by "{text}". We adhere to [22] to report $r=\textrm{max}(0,1-\frac{N_{e}}{N_{\textrm{ref}}})$, where $N_{e}$ is the minimum edit distance between the rendered text and the target text, and $N_{\textrm{ref}}$ is the non-whitespace length of the target text.
(3) Human preference alignment. We report the model-based reward outputs from Pickscore [30], HPSv2 [75], ImageReward [78] and UnifiedReward [72], which are models trained with large-scale human annotated preference data. We use the test prompts from DrawBench [56] to generate the images for evaluation.
(4) Image quality evaluation. We additionally report the CLIP-Score [53] and Aesthetic Score [57] on DrawBench [56], to evaluate the overall quality of generated images independent of the above task-specific criteria.
5.3 Results
Quantitative results. The evaluation results are shown in Tab. 1, where it can be seen that applying our proposed ARC on the T2I flow-matching model SD3.5-M for post-training yields consistent gains across task-specific, image quality, and human preference metrics. While improvements in human preference is modest, we observe a considerable improvement in performance on compositional generation (GenEval [21]), text rendering (OCR[14]) and CLIPScore [53], almost matching the performance of SD3.5-L in these three metrics, albeit having less than $\frac{1}{3}$ of the parameters (2.5B vs. 7.1B). This evidences that the text-to-image generative model’s intrinsic confidence is strongly correlated with the success of compositional generation, text rendering and text-image alignment, which are more objective metrics compared to human preferences.
Figure 4: Effect of ARC post-training SD3.5-M after post-training on PickScore [30] using FlowGRPO [37]. ARC complements external rewards, showing the best best compositional generation and visual appeal on GenEval [21]. Post-training on external rewards yields high visual appeal, but sacrifices compositionality as shown above (Column 3: Generates yellow motorcycle instead / generates unwanted human).
Figure 5: Qualitative results of ARC when applied on SD3.5 [16] on DrawBench [56], GenEval [21] and OCR [14]. It can be seen that applying ARC shows consistent improvements over the baseline SD3.5.
Task-specific Image Quality Human Preference
GenEval OCR ClipScore Aesthetic PickScore HPSv2.1 ImageReward UnifiedReward
Number of noise probes $K$
$K=4$ 0.71 0.66 0.287 5.37 22.34 0.273 0.81 3.08
$K=8$ (Ours) 0.71 0.67 0.288 5.39 22.41 0.278 0.87 3.11
$K=16$ 0.70 0.67 0.288 5.42 22.34 0.278 0.86 3.09
Classifier-Free Guidance for self-confidence calculation
O 0.68 0.59 0.287 5.38 22.39 0.278 0.85 3.10
X (Ours) 0.71 0.67 0.288 5.39 22.41 0.278 0.87 3.11
Offline vs Online Self-Confidence
Offline 0.69 0.61 0.285 5.36 22.36 0.274 0.82 3.07
Online (Ours) 0.71 0.67 0.288 5.39 22.41 0.278 0.87 3.11
Table 2: Ablation study results of ARC. We validate the design choices of ARC over number of noise probes $K$, the usage of CFG for self-confidence calculation, and online/offline self-confidence calculation. Our current configurations yield superior results.
We also analyze the effect of applying ARC post-training after post-training SD3.5-M on external rewards using Flow-GRPO [37]. The results show that while the performance on the external reward itself is mildly compromised, we consistently gain noticeable improvements across GenEval, OCR and CLIPSCcore, which strengthens our hypothesis that the intrinsic confidence of text-to-image generative model is strongly correlated with with the success of compositional generation, text rendering and text-image alignment. In Fig. 4, we provide visual examples of SD3.5-M post-trained with FlowGRPO using PickScore as the external reward, finally post-trained with ARC intrinsic self-certainty rewards. It shows that integrating ARC improves the generative performances even further, evidencing the complementarity of ARC and external rewards.
User study. In Fig. 3, we provide the results of a user study on prompts from PartiPrompt [81] and HPSv2 [75], asking users to assess the generated images based on visual appeal/realism and text alignment. We summarize $\sim$1,800 responses from 20 participants. The user study results show that applying ARC post-training SD3.5-M consistently outperforms the baseline model in terms of visual realism/appeal and text alignment.
Qualitative comparison. We provide additional qualitative comparisons in Fig. 1 and Fig. 5, where it can be seen that applying ARC yields visually pleasurable results with improved compositionality and text rendering capabilities, even when trained with no external reward at all.
5.4 Ablation study and analyses
In Tab. 2, we provide ablation study results to validate the design and hyperparameter choices of ARC.
Analyses on number of noise probes $K$. We vary $K$ across $4,8,16$ - the results show that while mostly similar, $K=8$ yields slightly better results overall. Although using $K=16$ outperforms $K=8$ in terms of aesthetic score, we decided that this improvement is negligible in comparison to the additional computation overhead during training.
CFG for self-confidence. We show that using CFG during self-confidence computation at post-training results in a slight drop in the overall performance. We conjecture this is because CFG itself is an inference-time technique, and using CFG inside the reward would optimize the guided proxy rather than the base conditional policy $\pi_{\theta}(\cdot\mid z_{t},c)$. This may incentivize reward hacking via larger guidance strength (larger than 1.0) instead of learning a better $\pi_{\theta}$.
Online self-confidence vs Offline self-confidence. We compare the post-training performance when the intrinsic reward is computed in an online manner i.e., calculated using the text-to-image model being trained $\pi_{\theta}$, and when computed in an offline manner i.e., calculated using the base model $\pi_{\textrm{ref}}$. The results show that using offline self-certainty as a static reward results in a performance drop across performance metrics, hinting that static self-confidence reward is suboptimal in comparison to using online self-certainty.
Observed causes of training collapse. We observe that the training collapses when (1) we train on a higher number of sampling timesteps i.e., $\rho>0.6$ in $\lvert\mathcal{T}_{\mathrm{train}}\rvert=\lceil\rho\,\lvert\mathcal{T}\rvert\rceil$, and (2) when we do not use CFG for sampling the $G$ candidates. We observe that under the two aforementioned cases, over-optimization against the intrinsic self-confidence reward occurs, resulting in textureless images being generated due to severe reward hacking and train collapse. We guide the authors to the supplementary for more details.
5.5 Limitations
One limitation of ARC is that the intrinsic self-confidence does not align strongly with human preference, where observed gains are modest. Also, while ARC shows to improve compositional generation, text rendering and text faithfulness, we cannot target a specific alignment objective using ARC alone. However, we showed that ARC can be integrated with external rewards to target specific alignments while alleviating reward hacking and improving compositionality of text rendering capabilities (Tab. 1).
6 Conclusion
We introduced ARC, a reinforcement-learning post-training framework that replaces external critics with an intrinsic self-confidence signals. ARC converts this self-confidence into stable scalar rewards and optimizes a flow-matching text-to-image generator without auxiliary annotators or task-specific validators. Across benchmarks and a user study, ARC yields consistent improvements in compositionality, text rendering and text alignment; a comprehensive user study supports our findings. We also showed that ARC can be integrated with external rewards to align text-to-image generation to specific external criteria with mitigated reward hacking and additional improvements. Future directions include (i) consistency-aware extensions (temporal/multi-view) to carry ARC to video and 3D generation, and (ii) disentangling and calibrating intrinsic signals to enable precise, task-targeted reward shaping.
[…]
Supplementary Material
Supplementary Contents
7 Rationale of ARC
We motivate ARC by testing whether a model’s denoising-based self-confidence correlates with perceived image quality when generation settings are made progressively stronger while keeping the underlying model fixed. Specifically, we compare three inference regimes for the same text–to–image model: (i) $10$ sampling steps without CFG, (ii) $10$ steps with CFG, and (iii) $20$ steps with CFG in Fig. 6. For each regime, we compute the self-confidence score defined in Sec. 4.1 on the model’s own outputs (same scorer and parameters across all cases). Empirically, the self-confidence distribution increases from (i) to (ii) to (iii), matching the observed rise in visual fidelity and prompt adherence. Because the signal is computed post hoc by the same model regardless of guidance or step count, these trends are not artifacts of an external evaluator; rather, they indicate that better samples are also easier for the model to self-denoise. This observation provides the rationale for treating denoising-based self-confidence as a model-native reward for post-training.
Figure 6: Rationale of ARC. Distributions of the denoising-based self-confidence under three inference settings—$10$ steps (no CFG), $10$ steps (CFG), and 2020 steps (CFG). The distribution shifts monotonically rightward (higher self-confidence) in the same order that visual quality improves, indicating that the ability to recover injected noise is predictive of sample quality even when the scorer is the same model. This alignment underpins ARC’s use of self-confidence as an intrinsic reward.
8 ARC Post-Training on SD3.5-L
To assess scalability, we apply ARC to SD3.5-L [16], a larger base model than the SD3.5-M used in the main experiments. Unless otherwise noted, we reuse the same training recipe (shortened denoising horizon, suffix-only updates, shared probes, CFG-free scoring). As reported in Tab. 3, ARC yields consistent gains in compositional generation, text rendering, and text–image alignment, while remaining competitive on human-preference metrics (e.g., HPSv2, PickScore). These results suggest that ARC scales to higher-capacity text-to-image models without inducing reward hacking and remains effective beyond the SD3.5-M setting.
9 Applying ARC on FLUX.1-Dev
To test architectural generality, we apply ARC to FLUX.1 - Dev [3], a flow-matching text-to-image generator with a design distinct from SD3.5. We keep the core ARC recipe unchanged (shortened denoising horizon, suffix-only updates, shared probes, CFG-free scoring), adapting only to the model’s native scheduler and inference step count. A small deviation is the suffix window: we set $\rho=0.5$, i.e., train on the latter half of the scheduler steps, which increased training stability in this setting. As reported in Tab. 3, ARC delivers consistent gains in compositional generation, text rendering, and text–image alignment, while remaining competitive on human-preference metrics (e.g., HPSv2, PickScore). The results indicate that ARC transfers effectively across architectures and remains robust on another representative flow-matching T2I model.
Task-specific Image Quality Human Preference
GenEval OCR ClipScore Aesthetic PickScore HPSv2.1 ImageReward UnifiedReward
SD3.5-M (2.5B) 0.65 0.61 0.282 5.36 22.34 0.279 0.84 3.08
+ ARC (Ours) 0.71 0.67 0.288 5.39 22.41 0.278 0.87 3.11
[…]
(Reproduced) 0.51 0.68 0.284 5.28 21.86 0.264 0.70 2.98
+ ARC (Ours) 0.58 0.74 0.288 5.25 21.91 0.253 0.65 2.98
[…]
(Reproduced) 0.66 0.61 0.269 5.71 22.84 0.274 0.88 3.21
+ ARC (Ours) 0.66 0.65 0.271 5.67 22.69 0.292 0.90 3.23
Table 3: Applying ARC to SD3.5-L [16] and FLUX.1-Dev [3]. We apply ARC on additional models of SD3.5-L and FLUX.1-Dev, to verify the effect of ARC given (1) a larger base model, and (2) a different architecture from SD3.5-M. $\dagger$ denotes results taken from DiffusionNFT [91]. We base our experiments on our reproduced results based on the official weights of SD3.5-L [16] and FLUX.1-Dev [3]. The results show that ARC consistently results in improved compositionality, text rendering and text-image alignment, while being competitive at human preference metrics.
10 Training Collapse Analysis
When and why collapse occurs. We monitor the batch-mean self-confidence (negative log error, averaged over probes and probed timesteps) across training iterations. Collapse is characterized by a rapid, sustained surge in this score—an overconfidence spike—followed by degenerate, low-texture generations (reward hacking). Empirically, two settings precipitate this behavior: (i) training on too many timesteps ($\rho>0.6$ in $|\mathcal{T}_{\mathrm{train}}|=\lceil\rho\,|\mathcal{T}|\rceil$), which exposes early, easily exploitable steps; and (ii) sampling the $G$ rollout candidates without CFG, which reduces exploration and inflates apparent self-confidence. A KL anchor alone is insufficient to prevent these modes.
Mitigations used in ARC. We restrict training to the latter $60\%$ of steps ($\rho=0.6$), keep CFG on during rollouts (but off when scoring self-confidence), and retain clipping, per-timestep weighting, and antithetic probes. These choices suppress overconfidence spikes and stabilize learning.
Figure 7: Visualization of training collapse in ARC. Self-confidence (y-axis) versus training iteration under different settings. Using $\rho>0.6$ or sampling rollouts without CFG drives a steep, short-horizon increase in self-confidence, followed by degenerate outputs—evidence of reward hacking. ARC’s default settings ($\rho{=}0.6$ and CFG for rollouts) avoid this behavior while preserving steady improvements.
11 Additional Ablation Studies
We conduct additional ablation studies and comparative experiments to validate the design choices of ARC. The results are summarized in Tab. 5.
11.1 Caption datasets for ARC
ARC relies on intrinsic self-confidence and thus requires only prompts (not external reward models). We compare three prompt sources: (i) text-rendering (OCR) prompts from Flow-GRPO [37] (our default), (ii) PickScore [30] prompts, and (iii) GenEval [21] prompts. As shown in Tab. 3, denser, more prescriptive prompts (OCR) yield the strongest gains; empirically, self-confidence is most reliable when the text condition is explicit and descriptive. We provide the descriptions and examples for each prompt dataset in Tab. 4.
(i) Text-rendering (OCR) prompts — default
Characteristics Dense, explicit textual content (exact strings, font/placement hints), strong conditioning for legibility and alignment.
Examples “A postage stamp design featuring the motto ”Unity in Diversity”, showcasing a vibrant collage of people from various ethnic backgrounds, each holding hands in a circle, set against a backdrop of colorful, interwoven patterns symbolizing unity and cultural richness.”
“In a luxurious hotel lobby, an elegant digital display above the elevator reads ”Now Playing”. Soft, ambient elevator music fills the space, enhancing the serene and welcoming atmosphere. A plush, modern sofa and a glass coffee table are seen in the foreground, with polished marble floors reflecting the ambient light.”
“A sleek, modern corporate lobby featuring a large, minimalist sculpture prominently inscribed with ”Innovate or Perish”, reflecting the company’s commitment to forward-thinking. The sculpture stands against a backdrop of glass and steel, with subtle lighting enhancing its form and the powerful message it conveys.”
(ii) PickScore prompts
Characteristics Open-ended statements; often adds context with simple concatentation of adjectives; weaker constraints on text content/layout.
Examples “An attractive young woman petting a cat”
“(a girl in steampunk fantasy world), (ultra detailed prosthetic arm and leg), (beautifully drawn face:1.2), blueprints, (magic potions:1.4), mechanical tools, plants, (a small cat:1.1), silver hair, (full body:1.2), magic dust, books BREAK (complex ultra detailed of medieval fantasy city), (steampunk fantasy:1.2), indoors, workshop, (Steam-powered machines:1.2), (clockwork automatons:1.2), (a small wooden toy), (intricate details:1.6), lamps, colorful details, iridescent colors, BREAK illustration, ((masterpiece:1.2, best quality)), 4k, ultra detailed, solo, (photorealistic:1.2), asymmetry, looking at viewer, smile”
“Cyborg cow, cyberpunk alien india, body painting, bull, star wars design, third eye, mehendi body art, yantra, cyberpunk mask, baroque style, dark fantasy, kathakali characters, high tech, detailed, spotlight, shadow color, high contrast, cyberpunk city, neon light, colorful, bright, high tech, high contrast, synthesized body, hyper realistic, 8k, epic ambient light, octane rendering, kathakali, soft ambient light, HD,”
(iii) GenEval prompts
Characteristics Compositional verification (objects, counts, relations), moderate specificity, minimal typography.
Examples “a photo of a yellow bus and an orange handbag”
“a photo of four surfboards”
“a photo of a book left of a cat”
Table 4: Prompt sources compared for ARC. Denser, text-focused prompts (OCR) provide stronger supervision signals for intrinsic self-confidence, leading to larger gains than more open-ended (PickScore) or simple compositional (GenEval) prompts.
11.2 Effect of group size
We clarify a typographical error in the main paper: although we stated G=24G{=}24, all experiments used G=16G{=}16. Varying $G$ shows that G=16G{=}16 outperforms G=8G{=}8 (more within-prompt exploration improves group-relative normalization) while G=32G{=}32 destabilizes training: larger groups reduce the number of distinct prompts per batch, lowering inter-prompt diversity and increasing the risk of over-optimization under relative advantages. In practice, G=16G{=}16 strikes a robust compute–stability trade-off.
11.3 Stepwise vs. aggregated reward
Although ARC’s self-confidence can be computed per step, we find that using the aggregated reward, i.e., averaging weighted per-step scores over the probed timesteps, consistently performs better than optimizing stepwise advantages. Stepwise improvements at individual timesteps need not translate to a better final sample and tend to increase variance and solver sensitivity; aggregation provides a more stable, outcome-aligned signal for post-training.
Task-specific Image Quality Human Preference
GenEval OCR ClipScore Aesthetic PickScore HPSv2.1 ImageReward UnifiedReward
Caption dataset used for ARC
- 0.65 0.61 0.282 5.36 22.34 0.279 0.84 3.08
PickScore prompts 0.70 0.62 0.285 5.26 22.13 0.278 0.65 2.96
GenEval prompts 0.71 0.62 0.286 5.32 22.35 0.275 0.80 3.05
OCR prompts (Ours) 0.71 0.67 0.288 5.39 22.41 0.278 0.87 3.11
Group size $G$
8 0.70 0.64 0.285 5.29 22.28 0.267 0.75 3.00
16 (Ours) 0.71 0.67 0.288 5.39 22.41 0.278 0.87 3.11
32 0.61 0.51 0.274 5.18 21.73 0.226 0.16 2.73
Step-wise reward vs. Aggregated reward
Stepwise 0.67 0.60 0.285 5.39 22.36 0.277 0.83 3.07
Aggregated (Ours) 0.71 0.67 0.288 5.39 22.41 0.278 0.87 3.11
Table 5: Additional ablation/comparative results. The results show that our current design choices for the (1) Caption dataset used, (2) Group size $G$, and (3) Aggregated self-confidence rewrads yield the best performances.
12 Additional Implementation Details
In this section we summarize the main implementation choices used in our ARC training pipeline. We acknowledge and correct a typographical error in the main paper: although we stated that the group size was $G=24$, all experiments were in fact conducted with $G=16$. The summary of hyperparameterse and configurations is illustrated in Tab. 6.
Category Hyperparameter Value (ARC, SD3.5-M)
Model Base model stabilityai/stable-diffusion-3.5-medium (SD3.5-M)
Components trained Transformer (denoiser) only; VAE and all text encoders frozen
LoRA LoRA usage use_lora = True
Rank $r$ $32$
Scaling factor $\alpha$ $64$
Init of LoRA weights Gaussian
Target modules attn.add_k_proj, attn.add_q_proj,
attn.add_v_proj, attn.to_add_out,
attn.to_k, attn.to_q, attn.to_v, attn.to_out.0
Data / prompts Train / test files train.txt, test.txt (one prompt per line)
Tokenization SD3.5 tokenizers; max length 128 (embeddings), 256 (logging)
Sampling Image resolution $512\times 512$
Sampler steps (train / eval) train:10, eval:40
Train timestep fraction train.timestep_fraction = 0.99 $\Rightarrow T_{\mathrm{train}}=9$
Suffix proportion $\rho$ in GRPO 0.6
Guidance scale (train/eval) sample.guidance_scale = 4.5
Noise level (SDE step) sample.noise_level = 0.7
Train batch size / GPU (sampling) sample.train_batch_size = 8 images
Test batch size / GPU sample.test_batch_size = 16 images
Images per prompt (group size $G$) sample.num_image_per_prompt = 16
Number of GPUs 88
Batches per epoch (sampling) sample.num_batches_per_epoch = 4
Global samples / batch 88 (bs) × $\times\,8$ (GPUs) =64=64 images
Prompts / batch $64/16=4$ prompts per sampling batch
Same latent per prompt sample.same_latent = False
Self-confidence (ARC) Probes per step $K$ 88 (antithetic pairing: $K/2$ noise, $K/2$ negated)
Probe timesteps Last half of used timesteps: $j=4,\dots,8$ (for $T_{\mathrm{train}}=9$)
Noise schedule for probe $\lambda_{t}=\tau_{t}/1000$; $x_{t}=(1-\lambda_{t})x_{0}+\lambda_{t}\epsilon$
Per-step score $s_{t}=-\log(\mathrm{MSE}_{t}+10^{-6})$, MSE between injected and predicted noise
Normalization Per-timestep batch-wise z-score, then mean over timesteps
CFG inside probe Disabled (conditional branch only)
[…]
KL regularizer weight train.beta = 0.04
KL form $D_{\mathrm{KL}}=\|\mu_{\theta}-\mu_{\mathrm{ref}}\|_{2}^{2}\big/(2\sigma_{t}^{2})$ (mean-only Gaussian)
Optimization / EMA Optimizer AdamW on LoRA parameters (no base-parameter updates)
Learning rate $3\times 10^{-4}$ (constant)
Gradient clipping Global norm clipping at train.max_grad_norm
EMA usage train.ema = True
EMA decay $0.9$
EMA update interval Every $8$ optimizer steps (update_step_interval = 8)
EMA usage in eval EMA weights used for evaluation; online weights restored afterwards
External rewards / eval Training reward Internal self-confidence only (no external reward in training)
SDS-only eval Optional SDS self-confidence evaluation on EMA model for monitoring
Table 6: Hyperparameters and key implementation details for ARC training on SD3.5-M.
12.1 Base models and LoRA configuration
We build on the StableDiffusion3Pipeline from diffusers with the pretrained model SD3.5-M: stabilityai/stable-diffusion-3.5-medium. We freeze all components except the denoiser: the VAE and all text encoders are kept fixed and used only for inference. Only the main transformer (UNet-like denoiser) is updated during training, based on LoRA. We run the text encoders in mixed precision (fp16 in our main ARC runs) and keep the VAE in fp32 for stability.
For parameter-efficient fine-tuning we apply LoRA to the transformer with
  • LoRA rank $r=32$ and scaling factor $\alpha=64$,
  • Gaussian initialization of LoRA weights,
  • Target modules inside each attention block:
    attn.add_k_proj, attn.add_q_proj, attn.add_v_proj, attn.to_add_out,
    attn.to_k, attn.to_q, attn.to_v, attn.to_out.0.
All non-LoRA base weights remain frozen.
12.2 Datasets and prompt processing
We consider two kinds of prompt datasets:
  • Plain text prompt datasets. We store the prompts in plain text files train.txt and test.txt. Each line contains a single prompt string. (e.g. PickScore, Text Rendering dataset)
  • GenEval-style metadata. For experiments on GenEval-style prompts we use JSONL files {train,test}_metadata.jsonl, where each line is a JSON object that contains at least a "prompt" field and additional metadata.
For each batch of prompts we compute text embeddings using the three SD3.5 text encoders. We also precompute embeddings for the empty prompt "" and use them as unconditional embeddings for classifier-free guidance (CFG) during sampling and log-probability computation.
12.3 Distributed sampling and grouping
[…]
12.4 KL regularization
Following Flow-GRPO, regularize the policy via a KL term that constrains the transition mean to stay close to a reference (the base model without LoRA):
  • The SDE step module returns the current mean $\mu_{\theta}$ and a reference variance $\sigma_{t}^{2}$.
  • We compute a reference mean $\mu_{\text{ref}}$ by temporarily disabling LoRA adapters and re-evaluating the same step.
  • Assuming Gaussian transitions with equal variance, the per-step KL divergence simplifies to
    $D_{\mathrm{KL}}=\frac{1}{2\sigma_{t}^{2}}\left\|\mu_{\theta}-\mu_{\text{ref}}\right\|_{2}^{2}.$
We average this KL over spatial dimensions and the batch and add it to the policy loss with weight $\beta=0.04$.
13 User Study Instructions and Interface
We provide the details of the instructions and interface used for the user study.
Instructions.
For each text prompt, you will be shown a pair of AI-generated images (left and right). For every image pair, you are asked to answer the following two questions independently:
  1. 1.
    Visual realism and appeal: Which image do you find to be more visually realistic and appealing?
  2. 2.
    Text–image alignment: Which image better aligns with the given text description?
For each question, please select your preferred image (left or right) based solely on the specified criterion.
Interface.
The user interface used in the study is illustrated in Fig. 8.
Figure 8: User study interface used to collect human preferences between pairs of AI-generated images.
14 Additional Qualitative Results
We provide side-by-side samples for (i) PickScore–post-trained (Flow-GRPO) SD3.5–M, (ii) FLUX.1–Dev, and (iii) SD3.5–L in Fig. 9. Across diverse prompts, ARC yields visibly sharper text rendering, more faithful object counts and relations, and fewer artifacts, echoing the quantitative gains in compositionality, text rendering, and text–image alignment, with no obvious regressions on non-target aspects.
Figure 9: Additional qualitative results of ARC. We present additional qualitative results of ARC when applied to (1) Flow-GRPO [37] post-trained SD3.5-M [16], (2) FLUX.1-Dev [3], and (3) SD3.5-L [16]. Best viewed on electronics.
Acknowledgement. This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2025-02653113, High-Performance Research AI Computing Infrastructure Support at the 2 PFLOPS Scale)